OpenStudy (anonymous):

ok so can someone help me with this question -3/4(-x+8)=1/2(6x-10) i know about the first step of distribution property but thats about it

OpenStudy (anonymous):

@animal_lover36

OpenStudy (anonymous):

do you need to solve for x?

OpenStudy (anonymous):

is the question like this : \(\Large \frac{-3}{4(-x+8)}=\frac{1}{2(6x-10)}\) or \(\Large\frac{-3}{4}(-x+8)=\frac{1}{2}(6x-10)\) ?

OpenStudy (anonymous):

to answer your first question yes but the second one is no the problem only has 2 fraction which is a negative 3 4ths and a 1 half

OpenStudy (anonymous):

@Data_LG2 hey sorrry i was helping my mom with something

OpenStudy (anonymous):

no problem :) so it's like this : \(\Large\frac{-3}{4}(-x+8)=\frac{1}{2}(6x-10)\)?

OpenStudy (anonymous):

Yes thats exactly it except the negative is infront of the actual fraction bar not infront of the 3

OpenStudy (anonymous):

dont worry it's the same thing :) so first, let's get rid off the fractions. we can do this by cross multiplying the denominator